Imagine that for the unimolecular reaction A LaTeX: \Longleftrightarrow⟺ B, the value of LaTeX: \DeltaΔG° is -1.17 kcal/mol. If equimolar initial amounts of A and B react at 25 °C, what fraction of the molecules will be B molecules once the system has come to equilibrium?
Answer #1

ΔG° = -1.17 kcal/mol

ΔG° = - R T ln K

- 1.17 x 10^3 = - 1.987 x 298 x ln K

Keq = 7.214

fraction of B molecules = Keq / Keq + 1

                                     = 7.214 / 7.214 + 1

fraction of B = 0.88

                       = 88 %